OpenStudy (anonymous):

Find the slope of the line with these given points (9, 3) and (3, -6). A. 3/2 B. 2 C. 1/3 D. -1/3

OpenStudy (anonymous):

(9-3)/(3-(-6))=6/9=2/3

OpenStudy (anonymous):

\[\huge \color{green}{Slope = \frac{y_2 - y_1}{x_2 - x_1}}\]

OpenStudy (anonymous):

@marktierra but 2/3 isnt one of the answers

OpenStudy (australopithecus):

use waterineyes formula

OpenStudy (australopithecus):

(x,y) (9,3) (3,-6)

OpenStudy (anonymous):

\[Slope= \frac{-6-3}{3-9} \implies \frac{-9}{-6} \implies \frac{3}{2}\]

OpenStudy (anonymous):

thanksX)

OpenStudy (anonymous):

@marktierra , check your solution according to the formula..

OpenStudy (anonymous):

Welcome dear..